Look at the circuit below j312 - 3412 Ro = 22 VO 420° A 620° V + + 412Removing the 2 ohm load resistor and looking into the terminals where Vo is measured, what is the Thevenin equivalent impedance of the circuit in rectangular coordinates
Answer #1


Thevenin Impedence: (Zth) Thevenin. impedence is the voltage source impedence across 22 with shorted and current source opene(4-4) 2 and jar are connected in paollel : Zth = (4-34) paollel to j3r .: Zah (4-54) C j3) 4-34 +j3 : Zih S -34(j3) 33(4) j 1-) Zuh = 36 +j6o 17 Zth = い 36 17 tj bo 17 Zth = (2.1176 +j 3.5294)2 -) The venin equivalent impedence Zth (2.11 2.11767; 3.5